2 2/3 divided by 1 3/5
1
Expert's answer
2012-09-21T10:23:50-0400
First we should transform these mixed numbers into fractions:
2 2/3=(6+2)/3=8/3
1 3/5=(5+3)/5=8/5
Now let's divide:
8/3/(8/5)=8*5/(8*3)=5/3